High Performance Computing Service

Relocation Schedule

Last update:Sat 28 Feb 17:01:36 GMT 2015

All HPCS services will be moving to the West Cambridge Data Centre (WCDC) during January and February 2015. The current plan is outlined on this page and any amendments will be reflected here.

When  Event  Comments
  
Friday 23rd January 08:00  Temporary reduction in capacity.
DONE
  Withdrawal of 60% of Teslas, and reduction of the number of available Sandy Bridge nodes by 50%. Login and filesystem access, and job scheduling to the remaining resources, will continue uninterrupted. Hosted clusters will not be affected at this stage.

Phase 1 relocation is 100% complete.
  
Thursday 29th January 08:00   All production services suspended.
DONE
  No service - total blackout including login, filesystems and hosted clusters. This period will continue until at least the core infrastructure has been brought up in the WCDC and tested.

Phase 2 relocation is now underway - no access to clusters.
Physical moves 100% complete. The Mill Lane Data Centre now stands empty:


The Mill Lane Data Centre stands empty
  
Monday 9th February  Login and filesystem access to be restored when possible.  The core infrastructure (part of which can be seen below) is now powered up and all three main networks are built and functioning. However there is a thermal loading problem in the room affecting the racks containing storage and compute nodes which will require engineering attention on Monday. When this is resolved we expect to be able to reinstate partial service.

Core Infrastructure
Lustre filesystems are now ready to mount and the core is ready to provision some cluster nodes.
  
Tuesday 10th February      Login access resumed at 11:30.
  
Wednesday 11th February      SLURM, remote visualisation node and tape backups now all working. More than 50% production nodes now up with some issues. Hosted clusters nearing completion. Cooling system work in progress.
  
Saturday 14th February      Partial service on the Sandy Bridge nodes has been restored. At the moment no more can be opened up as a result of the incomplete cooling system work.  
  
Monday 16th February      Partial service on the Tesla nodes has been restored.  
  
Friday 27th February      Following further progress with the installation of additional cooling doors, all Sandy Bridge racks are now running jobs. 13 Tesla racks are now also running jobs.  
  
Saturday 28th February      All Sandy Bridge racks are 14/16 Tesla racks are now also running jobs.